OpenStudy (anonymous):

What is the area of a circle with circumference of 9.42 inches? (Use 3.14 for p.)

OpenStudy (cggurumanjunath):

a=pi*r^2 c=2*pi*r=9.42

OpenStudy (cggurumanjunath):

pi*r=9.42/2

OpenStudy (cggurumanjunath):

r=9.42/(2*pi)

OpenStudy (cggurumanjunath):

substitute above r in a=pi*r^2

OpenStudy (anonymous):

its 9.42